OpenStudy (anonymous):

Write the equation in standard form. Identify the important features of the graph: x^2 + y^2 - 8x + 10y + 15 = 0

OpenStudy (anonymous):

in the above 2 degree equation the term containing xy is absent and morever the coefficient of x^2 and y^2 are each equal to 1 hence the above equation may represent a circle

OpenStudy (anonymous):

x^2 +^2 - 8x + 10y + 15 = 0 try making perfect squares to bring it in the standard form

OpenStudy (anonymous):

x^2 -2(4)(x) +16+ y^2 + 2(5)y +25-25+ 15-16 = 0 (x-4)^2 + (y+5)^2 = 26 so the centre of the circle is at (4,-5) and radius is sqrt(26) units

OpenStudy (anonymous):

Compare your equation with the standard form, \[\LARGE x^2 + y^2 + 2gx + 2fy + c =0\] g=? f=? Centre of the circle: \((-g,-f)\) Radius:\(\Large \sqrt{g^2+f^2-c}\)

OpenStudy (anonymous):

are you solving sums on circles...

OpenStudy (anonymous):

Um I don't know....(I'm sorry a family member passed away and I missed like a month of school so I'm just completely lost)

OpenStudy (anonymous):

How would I solve the sums?